The line joining the points (-1,3) and (4,-2) will pass through the point (p,q) if

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

first find the equation of line passes through (-1,3) and (4,-2)

slope of line  m = (-2-3 )/(4+1)

                  =-1

equation of line  y-3 =-1(x+1)

                  y+x =2

 point(p,q)  lie on this equation

so  p+q =2

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin